Projection Admin 11/14
To do this evening/ was completed: We had Jack's review today. Had a good talk about all things projection related. Alec and I have good feedback for what we can work on department wise, too.
Between the Mountain and the Sky tech all good, volume and lights up logged (5.0, 1:30:00).
All A1 11/21 playlists built, plus two of the three SE's for the week.
No Other Choice posters in 3/7.
Presentation Issues:
Preshow Clips/Playlist:
Trailer Updates:
Special Events/ AV set up: Private event for Frankenstein in T3 went well today. Hosts showed a little late, but everything was perfectly fine. - L
Ordering/Receiving (posters & trailers):
Ordering/Receiving (content): Ingested Hellboy 2, Crow keys came in! (they open up this Sunday, so I will set a reminder, but most likely I'll take a look on Monday), DCP'd and ingested the Scott Barley Intro for Dark Room - L
Outgoing/Shipping (content):
Print/video Inspection: Went over the Dark Room shorts playlist and intro video with Crisinta in T3 today. Playlist updated with Scott intro. - L
Presentation Issues:
Preshow Clips/Playlist:
Trailer Updates:
Special Events/ AV set up: Private event for Frankenstein in T3 went well today. Hosts showed a little late, but everything was perfectly fine. - L
Ordering/Receiving (posters & trailers):
Ordering/Receiving (content): Ingested Hellboy 2, Crow keys came in! (they open up this Sunday, so I will set a reminder, but most likely I'll take a look on Monday), DCP'd and ingested the Scott Barley Intro for Dark Room - L
Outgoing/Shipping (content):
Print/video Inspection: Went over the Dark Room shorts playlist and intro video with Crisinta in T3 today. Playlist updated with Scott intro. - L
Ran Kiki's Delivery Service and Legends of the Fall for Cristina. Legends is Dolby A. I ran the Bugonia loop and adjusted the masking tape again for tonight's show. - L
Asset tracking doc updated/checked?:
Maintenance/cleaning: Replaced a woofer in the middle right side surround speaker in T4 -A
Asset tracking doc updated/checked?:
Maintenance/cleaning: Replaced a woofer in the middle right side surround speaker in T4 -A
Broke down the old bookshelf and took it out to the dumpster -A
DMS should be visiting next week to upgrade the T6 firmware, looking like Tuesday or Wednesday -A
Cue sheets updated/checked (every shift):
Marquee updated?:
Posters updated?: Nouvelle Vague is down and replaced with Jay Kelly. -A
Cleaning lights on for the night?:
Projection cart by entrance?:
Additional Notes:
Cue sheets updated/checked (every shift):
Marquee updated?:
Posters updated?: Nouvelle Vague is down and replaced with Jay Kelly. -A
Cleaning lights on for the night?:
Projection cart by entrance?:
Additional Notes:
Comments
Post a Comment